If D=100 what does N equal in the following equation? D = 3N + 1
Answer by Maths68(1474)

D = 3N + 1
D-1=3N
or we can write
3N=D-1
Divide by 3 both sides of above equation
3N/3=(D-1)/3
N=(D-1)/3
Put the value of D in above equation
N=(100-1)/3
N=99/3
N=33